OpenStudy (anonymous):

What is tan2Θ, if sinΘ = -24/25 and Π <Θ<3Π/2

OpenStudy (mertsj):



OpenStudy (mertsj):

\[\tan 2\theta =\frac{2\tan \theta}{1-\tan ^2\theta}\]

OpenStudy (anonymous):

What am I suppose to do after that ? @mertsj